# RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy in Language Modeling

### v1

@@ 3 Experiments
In this section, we present the efficiency (training, prefilling, and generation) and precision (pretraining, evaluation, and SFT) of different models. We focus primarily on the 1.3B model scale with a 4K context window, while additional discussions, including length generalization, 16K context window training, and more evaluation results, are presented in Appendix C.
@@ 3.1 Setup
For brevity, we summarize the setup of the 1.3B model with the 4K context window, which is used in most of our experiments, and provide full details in Appendix A.
Model.
We adopt a Transformer architecture that interleaves a token mixing block with a hidden-state mixing block (FFN), each wrapped with residual connections and LayerNorm. We compare variants using different token mixing modules, including RNN, attention, and RAT. Inspired by recent designs that alternate global and sliding-window attention (SWA), we also explore hybrid models. Specifically, Attention-SWA alternates standard attention with SWA (window size 1024) every two layers, while RAT-SWA replaces global attention with RAT.
@@
We benchmark the latency of a single token mixing block, including input and output projections, on the single H100 in the GH200 system with 120GB of memory. For training and prefilling, we report the latency of processing the entire sequence. For generation, which is often the most time-consuming case, we report the latency of generating a batch of tokens at a given position. To ensure a fair comparison with strong baselines, we use flash attention for the attention module and the associative scan operator for RNN. For RAT, we follow our method design as described above. All models are wrapped with torch.compile and evaluated in bfloat16 using torch.cuda.amp.
@@
We pretrain the 1.3B models on a 100B token subset of the FineWeb-Edu dataset [20]. We use a learning rate of 8.0e-4, decayed to 1.0e-6 using cosine annealing, and a global batch size of 2M tokens. Hyperparameters are determined by DeepSeek’s attention-optimized guidelines [21] and are shared across all model variants for fair comparison. Then, we further evaluate our models using two types of benchmarks. The first is a suite of classical commonsense reasoning tasks [22] which are commonly used to assess the reasoning ability of pre-trained models in short contexts. The second is a set of English tasks from LongBench [9], where the average input length for nearly every task exceeds 4,000 tokens, and some tasks have average lengths over 10,000 tokens. LongBench tasks often include specific prompts and instructions, which pretrained-only models struggle to follow reliably. To better assess the model capability, we also compare different architectures after SFT. Specifically, we choose NarrativeQA [23] with two modes, QMSum [24], and WikiSum [25], which vary in sequence length and remain close to the pretraining domain.
3.2 Efficiency study
Fig. 3 presents the efficiency comparison of RNN, RAT, and attention across three scenarios: training, prefilling, and generation. For training and prefilling, when the sequence length is short (e.g., 4096), RAT is slightly slower than attention due to under-utilized GPU parallelism in the flex attention component (with only 256 chunks) and the overhead of associative scan. We believe this can be improved with better CUDA kernel optimization. As the sequence length increases, RAT becomes more efficient: for $L=16$, we observe about 2×2$\times$2 × speedup at 16K, 3×3$\times$3 × at 32K, 4×4$\times$4 × at 64K, and 7×7$\times$7 × at 100K tokens. Additional results in Subsec. B.1 show up to 910×9-10$\times$9 - 10 × at 200K. While RNN slows down beyond 32K, RAT benefits from increased intra-chunk parallelism and retains an advantage over attention. For generations, RAT(L=16) achieves 9×9$\times$9 × speedup at position 4K and around 10×10$\times$10 × for longer sequences. It also reduces the usage of the KV cache, making it much less prone to OOM. RAT(L=64) can be up to 4×4$\times$4 × faster than $L=16$ with larger batch sizes or longer inputs.
Figure 3: Latency results. Latency of a single RNN, RAT, or attention layer (including input and output projections) with model dimension of 2048 under (a) training, (b) prefilling, and (c) generation. (a,b) report full-sequence latency on 200K tokens; (c) reports latency for generating 512 tokens at a fixed position. RAT also reduces KV cache usage, lowering OOM risk. See Appendix B for more results.
3.3 Ablation study
@@
3.4 Pretraining
We begin our study by examining pretraining perplexity as a foundational measure of architectural performance, showing that RAT lies between attention and RNN. The main results are presented in Fig. 4(b) and (c). In (b), on a 200M model, increasing the size of the chunk leads to higher perplexity. Viewing attention as $L=1$ and RNN as $L=T$, we observe perplexities of 13.26 for attention, 13.42 for RAT(L=64), 13.72 for RAT(L=256), and 14.31 for RNN. Interestingly, increasing the training context from $T=8192$ to $T=32768$ causes a sharp increase in perplexity for RNN, while attention and RAT remain stable. Note that on a 200M model, the PPL may not decrease steadily when going to longer sequences. We then scale up to a 1.3B model trained on web data, which primarily consists of short sequences (often under 2K tokens) concatenated with separator tokens to construct 4K-length contexts. In this regime, we find that the chunk length should be reduced to 16 to match the performance of attention. Once again, increasing the size of the chunks and thus reducing the FLOPs leads to higher perplexity, strengthening the trade-off and intermediate nature of RAT.
3.5 Downstream evaluation
Table 2: We evaluate our model on six commonsense reasoning tasks using the lm-evaluation-harness [22]. These tasks have very short contexts, with each sample containing no more than 300 tokens. Normalized accuracy is reported when available, except for WinoGrande and LAMBADA, which use standard accuracy.
Model ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande Avg.
RNN 38.57 66.46 55.95 39.74 72.69 53.99 54.57
RAT(L=16) 39.33 67.80 56.44 44.40 72.69 53.20 55.64
Attention 37.71 66.67 57.44 47.84 72.91 57.70 56.71
Attention-SWA 37.20 66.20 57.09 47.97 72.74 56.91 56.35
RAT(L=16)-SWA 40.78 68.22 57.85 49.33 73.94 56.91 57.84
@@ 3.5 Downstream evaluation
The performance gap between RNN and attention is small on short and relatively simple commonsense reasoning tasks, typically within two points on average. This suggests that RNN can be surprisingly competitive in short-context settings. RAT(L=16) consistently outperforms RNN in most tasks, with particularly strong gains on LAMBADA, highlighting its advantage in modeling token dependencies. However, it performs poorly on WinoGrande compared to attention. One possible explanation is that WinoGrande relies heavily on subtle sentence-level cues and fine-grained referential resolution; moreover, the variance may also stem from the instability of the task itself. Interestingly, interleaving RAT with SWA yields a hybrid model (RAT(L=16)-SWA) that achieves the best overall performance in almost all tasks. This suggests a complementary architecture: using SWA for strong local attention and RAT for long-range dependencies can bring us a model that is not only more efficient but also more accurate than its dense counterpart. In addition, the hybrid version mitigates the weaknesses of RAT in WinoGrande, improving overall robustness. We do not include results from Gated DeltaNet [26], a recent state space model that also interleaves with sliding-window attention, [^1] While our best model outperforms their best model (e.g., 49.33 vs. 48.76 on LAMBADA, 73.94 vs. 72.57 on PIQA, and 57.85 vs. 56.88 on HellaSwag), we do not include their results in our tables. Their approach follows a fundamentally different modeling paradigm—state space modeling—whereas our work focuses on an intermediate design between recurrence and attention. These differing motivations, together with discrepancies in window size (their 2048 vs. our 1024) and training protocols, make direct comparisons less meaningful. as their modeling paradigm and experimental setup differ substantially from ours.
Table 3: Evaluation results on LongBench. For each task, we highlight the top two values to illustrate general trends on instruction-based tasks. To save space, task numbering follows the original LongBench paper; the number-to-task mapping and token counts (measured with the LLaMA2 tokenizer) are provided in Subsec. C.4. Metrics: F1 for QA tasks, Rouge-L for summarization, and EditSum for code tasks. Specifically, accuracy is used for 4-1, F1 for 4-2, and Rouge-L for 4-3.
@@
Model Single-Document QA Multi-Document QA Summarization Few-shot learning Code Completion
1-1 1-2 1-3 Avg. 2-1 2-2 2-3 Avg. 3-1 3-2 3-3 Avg. 4-1 4-2 4-3 Avg. 6-1 6-2 Avg.
RNN 11.7 8.8 19.5 13.3 9.1 15.1 5.4 9.9 16.9 16.8 16.4 16.7 10.0 17.5 13.6 13.7 14.0 17.6 15.8
RAT(L=64) 13.9 13.0 27.0 18.0 14.9 16.3 4.3 11.8 18.9 17.7 19.9 18.8 17.5 19.2 15.2 17.3 16.2 19.7 18.0
RAT(L=16) 14.5 16.1 28.3 19.6 16.7 18.9 6.3 14.0 24.9 17.5 18.3 20.2 38.5 18.2 18.7 25.1 14.2 20.6 17.4
Attention 12.3 14.0 28.2 18.2 13.4 17.7 4.9 12.0 18.6 18.9 21.0 19.5 39.5 19.6 23.9 27.7 21.3 26.5 23.9
Attention-SWA 13.1 14.6 24.5 17.4 13.7 19.0 6.1 12.9 19.4 17.3 21.4 19.4 47.0 19.6 25.1 30.6 18.7 24.6 21.7
RAT(L=16)-SWA 12.7 15.6 28.2 18.8 14.2 18.1 7.4 13.2 20.1 18.6 19.8 19.5 33.5 18.4 25.6 25.8 26.3 30.1 28.2
@@
We present the results of the evaluation for 14 English tasks in LongBench in LABEL:tab:longbench_4k. It is expected that no single model consistently dominates across all tasks, since LongBench employs heavy instruction-style prompts and the tasks span diverse domains. Evaluating pretrained-only models without any fine-tuning on these instruction-based tasks remains challenging; even in the original paper, 7B pretrained-only models perform much worse than tuned ones. Therefore, we focus on general performance trends across tasks, particularly average scores by task category and the top-2 models for each task.
First, RAT(L=16) and its hybrid variant RAT(L=16)-SWA consistently rank among the top two models in nearly all task groups. This is notable given that RAT(L=16) lags behind attention in pretraining perplexity and performs slightly worse in short-context commonsense reasoning. We hypothesize that its advantage on LongBench stems from structural adaptation to very long sequences—for instance, achieving 24.9 on GovReport (3-1) (vs. 18.6 for the attention baseline) and 16.7 on HotpotQA (2-1) (vs. 13.4). Second, unlike the small gap ( 2 points) between RNN and attention on short-context tasks, the difference becomes much larger on LongBench, reaching up to 8 points on code completion. Finally, we note some outliers: RAT-based models underperform on TREC (4-1), a few-shot QA task evaluated via classification accuracy, for reasons not yet fully understood. In addition, RAT(L=16) performs poorly on code completion. However, its hybrid variant RAT(L=16)-SWA significantly improves in this setting, outperforming the second-best model by more than four points, consistent with earlier observations on commonsense reasoning tasks.
3.6 Supervised Fine-tuning
Since LongBench tasks are instruction-based and pretrained-only models often struggle to follow specific prompts, we further assess the performance of SFT on the QA and summarization tasks in the long-context setting, as shown in Table 4.
For NarrativeQA, where passages are extremely long and may not contain the answer within the truncated 4K-token context, we evaluate two settings: (1) summary-only, which uses only the passage summary; and (2) summary+passage, which concatenates the summary and full passage. While the latter provides more context, it also introduces much noise. In both cases, RAT(L=16)-SWA delivers strong performance while remaining efficient. Although the summary and passage contain useful information, we observe that models, particularly the attention module, perform worse in the second setting, with F1 dropping to 33. This highlights the difficulty in identifying and focusing on relevant content in noisy, extended contexts. In this setup, attention can be more easily distracted due to the absence of structural bias, which can lead to flattened attention score distributions [27]. In contrast, models based on RAT appear to be more robust, likely due to their chunk-wise structure. For summarization tasks, RNN performs better than in QA, and RAT(L=16)-SWA achieves the best result on QMSum, surpassing the next-best model by approximately 1 Rouge-L point.
@@ 3.6 Supervised Fine-tuning
[…]

### v3

@@ 4 Experiments
In this section, we present the efficiency and large-scale evaluations of RAT, along with comparisons to other models. Additional discussions are provided in Appendix C.
@@ 4.1 Setup
For brevity, we summarize the setup for the 1.3B model with a 4K context window, which is used in most of our experiments. Full implementation details are available in Appendix A.
Model
We adopt a Transformer architecture that interleaves a token mixing block with a hidden-state mixing block (FFN), each wrapped with residual connections and LayerNorm. We compare variants that use different token mixing modules, including RNN, attention, and RAT, as well as recent state space and linear attention models. In addition, we provide hybrid models that interleave RAT with sliding-window attention (window size 1024), such as Attention-SWA and RAT-SWA.
@@
We benchmark the latency of a single token mixing block, including input and output projections, on a single H100 GPU (GH200 system, 120GB), and also report the maximum throughput of the full model. We measure the time required to train on a full sequence or to generate a batch of tokens at specific positions. For fair comparison, we use flash attention for the attention baseline and associative scan for RNN. All models are compiled using torch.compile and evaluated in bfloat16 with torch.cuda.amp.
@@
We pretrain all 1.3B models on 100B tokens from FineWeb-Edu [fineweb_edu] using the same setup: a learning rate of 8.0e-4 decayed to 1.0e-6 (cosine schedule) and a global batch size of 2M tokens, following DeepSeek’s hyperparameter guidelines [deepseek_hp]. We directly evaluate the models on both short-context tasks—several classical commonsense reasoning benchmarks [eval-harness]—and long-context tasks, including 11 tasks from LongBench [longbench] covering QA, summarization, and code completion. Since LongBench includes instruction-heavy prompts that pretrained models often struggle with, we also include SFT-based evaluations. Specifically, we use NarrativeQA [narrativeqa] (two modes), QMSum [qmsum], and WikiSum [wikisum] to test long-context understanding with SFT. To assess retrieval capabilities, we include the Ruler benchmark [ruler] and test nine synthetic needle-in-haystack tasks with varying configurations. A single round of lightweight fine-tuning is applied to adapt models to specific prompts.
4.2 Analyses of RAT
Figure 2: Latency of the temporal mixing block (including linear projections) with a model dimension of 2048. (a): full-sequence latency with 200K tokens; (b): generation of 512 tokens at specified positions. We adopt flash attention for Attention.
Table 2: Maximum throughput of full models (tokens/sec), measured by generating 1024 tokens from a 3072-token prompt. By reducing the KV cache memory and boosting speed, we achieve 10× maximum throughput compared to flash attention , and even more on 13B models, as attention suffers from poor GPU utilization at larger scale.
[…]
Efficiency study
Subsec. 4.2 and Table 2 present the efficiency comparison, with additional results provided in Appendix B. For training, on the strong H100 GPU, when the sequence length is short (e.g., 4096), RAT is slightly slower than attention due to underutilized GPU parallelism in the flex attention (with only 256 chunks) and the overhead introduced by associative scan. We expect these can be improved through further kernel-level optimizations. As the sequence length increases, RAT becomes increasingly efficient: for $L=16$, we observe approximately $2\times$ speedup at 16K, $3\times$ at 32K, $4\times$ at 64K, and $7\times$ at 100K tokens. For generation, RAT(L=16) achieves a $9\times$ speedup at position 4K and around $10\times$ for longer sequences. It also reduces KV cache usage, making it significantly less prone to out-of-memory (OOM) errors and enabling much higher maximum throughput.
Ablation study
@@ 4.2 Analyses of RAT
Pretraining
In Fig. 3(b) and (c), we begin our study by examining pretraining perplexity as a foundational measure of architectural performance, demonstrating that RAT falls between attention and RNN. In (b), using a 200M model, we observe that increasing the chunk size leads to higher perplexity while reducing FLOPs. Viewing attention as $L=1$ and RNN as $L=T$, we observe perplexities of 13.26 for attention, 13.42 for RAT(L=64), 13.72 for RAT(L=256), and 14.31 for RNN. Interestingly, increasing the training context from $T=8192$ to $T=32768$ leads to a sharp rise in perplexity for RNN, while both attention and RAT remain stable, demonstrating the pitfalls of full-sequence compression. We then scale up to a 1.3B model trained on web data, which primarily consists of short sequences (often under 1K tokens) concatenated with separator tokens to construct 4K-length contexts. We find that RAT(L=16) gives the comparable performance to attention, and again RAT with different chunk sizes exhibits behavior intermediate between attention and RNN.
4.3 Large-scale evaluation and comparisons
In this section, we present direct evaluation and SFT results of our trained 1.3B models. We mainly compare RAT(L=16) with its two extremes: RNN and attention. Additionally, we include results from recent state space and linear attention models [gateddeltanet], which are trained under the same settings (model size, sequence length, and dataset) as ours. We report the value/slot sizes, which indicates the size of memory slots and FLOPs, based on sequence length $T$, model dimension $D$, and number of layers $N$. Note that state space and linear attention models use fixed-size memory and expand the classic recurrence state, while RAT scales its memory capacity with the sequence length.
@@ 4.3 Large-scale evaluation and comparisons
In Table 3, the performance gap (within two points on average) between RNN and attention is small on these short commonsense reasoning tasks, indicating that RNN can be surprisingly competitive in short-context settings. RAT(L=16) consistently outperforms RNN on most tasks and surpasses other recent models on 4 out of 7 benchmarks, with only at most $20ND$ memory slots. We also observe that some binary-choice tasks (e.g., BoolQ, WinoGrande) show inconsistent results under small-scale pretraining. For instance, attention struggles on BoolQ, while RAT underperforms on WinoGrande. Interestingly, the hybrid variant RAT(L=16)-SWA, which interleaves RAT with SWA, achieves the best overall accuracy across almost all tasks. This suggests a complementary architecture: using SWA for strong local attention and RAT for long-range dependencies can bring us a model that is not only more efficient but also more accurate than its dense counterpart.
Table 3: We evaluate our models with model dimension $D=2048$ and number of layers $N=24$ on seven commonsense reasoning tasks using lm-evaluation-harness [eval-harness]. Since all tasks (except BoolQ, which has only 13 sequences >300>300 tokens) have sequences $\leq 300$ tokens, we set $T=300$. Results marked with * are from a recent study [gateddeltanet] trained on the same setting as ours. For SWA, slot sizes remain $150ND$ due to the short context.
Model Value/Slot size ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande BoolQ
$T=300$ acc/acc_n acc/acc_n acc/acc_n acc acc/acc_n acc acc
RNN NDND 35.15/38.57 70.66/66.46 42.43/55.95 39.74 71.49/72.69 53.99 62.54
Mamba* $64ND$ -/35.40 69.52/- -/52.91 -/43.98 71.32/- 52.95 61.13
Mamba2* $256ND$ -/37.88 72.47/- -/55.67 -/45.66 71.87/- 55.24 60.13
DeltaNet* $128ND$ -/35.66 68.47/- -/50.93 -/42.46 70.72/- 53.35 55.29
GatedDeltaNet* $288ND$ -/38.39 71.21/- -/55.76 -/46.65 72.25/- 57.45 60.24
Attention $300ND$ 35.67/37.71 71.25/66.67 44.16/57.44 47.84 73.29/72.91 57.70 58.23
RAT(L=16) $\sim 20ND$ 36.35/39.33 72.64/67.80 43.27/56.44 44.40 72.03/72.69 53.20 63.30
GatedDeltaNet-SWA* 288N/2D300N/2D288N/2D-300N/2D -/40.10 71.75/- -/56.53 47.73 72.57/- 58.4 63.21
Attention-SWA 300N/2D300N/2D300N/2D-300N/2D 35.24/37.20 71.97/66.20 44.36/57.09 47.97 72.25/72.74 56.91 61.25
RAT(L=16)-SWA $\sim 20N/2D-300N/2D$ 37.20/40.78 72.56/68.22 44.33/57.85 49.33 73.29/73.94 56.91 63.21
Table 4: Evaluation results on LongBench with $T=4096$, $D=2048$, and $N=24$. NQA: NarrativeQA, MQA: MultiFieldQA-en, HQA: HotpotQA, WQA: 2WikiMultihopQA, MSQ: Musique, GR: GovReport, MN: MultiNews, TQA: TriviaQA, RBP: RepoBench-P. Metrics: F1 for QA tasks, Rouge-L for summarization, and EditSum for code tasks. Results marked with * are from a recent study [gateddeltanet] trained under the same settings as ours. Note that they use local attention with a window size of 2048, while we use 1024.
@@
Model Value/Slot size Single-Document QA Multi-Document QA Summarization Code Completion
$T=4096$ NQA Qasper MQA Avg. HQA WQA MSQ Avg. GR QMSum MN Avg. LCC RBP Avg.
RNN NDND 11.7 8.8 19.5 13.3 9.1 15.1 5.4 9.9 16.9 16.8 16.4 16.7 14.0 17.6 15.8
Mamba2* $256ND$ 11.1 11.3 18.6 13.7 11.8 15.1 6.7 11.2 6.7 14.5 7.4 9.5 17.9 20.6 19.3
DeltaNet* $128ND$ 12.9 10.8 21.5 15.1 10.9 13.2 5.1 9.7 6.5 13.5 7.2 9.1 17.6 20.3 19.0
GatedDeltaNet* $288ND$ 14.1 14.0 23.3 17.1 13.7 14.4 5.8 11.3 7.5 16.4 7.9 10.6 18.7 22.1 20.4
Attention $4096ND$ 12.3 14.0 28.2 18.2 13.4 17.7 4.9 12.0 18.6 18.9 21.0 19.5 21.3 26.5 23.9
RAT(L=64) $64ND$ 13.9 13.0 27.0 18.0 14.9 16.3 4.3 11.8 18.9 17.7 19.9 18.8 16.2 19.7 18.0
RAT(L=16) $256ND$ 14.5 16.1 28.3 19.6 16.7 18.9 6.3 14.0 24.9 17.5 18.3 20.2 14.2 20.6 17.4
GatedDeltaNet-SWA* 288N/2D2048N/2D288N/2D-2048N/2D 14.5 12.3 26.6 17.8 12.6 23.6 6.1 14.1 9.1 16.1 12.8 12.7 15.5 19.2 17.4
Attention-SWA 4096N/2D1024N/2D4096N/2D-1024N/2D 13.1 14.6 24.5 17.4 13.7 19.0 6.1 12.9 19.4 17.3 21.4 19.4 18.7 24.6 21.7
RAT(L=16)-SWA 256N/2D1024N/2D256N/2D-1024N/2D 12.7 15.6 28.2 18.8 14.2 18.1 7.4 13.2 20.1 18.6 19.8 19.5 26.3 30.1 28.2
@@
In LABEL:tab:longbench_4k, we present results on 11 long-context tasks from LongBench. Since LongBench includes instruction-heavy prompts across diverse domains, it’s natural that no single model dominates all tasks. Evaluating purely pretrained models on such instruction-based datasets can be very challenging. Therefore, we focus on general performance trends, especially average scores by task category. As shown in the table, RAT(L=16) and its hybrid variant RAT(L=16)-SWA achieve top performance in many groups, such as question answering and summarization. This is noteworthy, given that RAT(L=16) slightly lags behind attention in pretraining perplexity and underperforms in short-context commonsense reasoning. We hypothesize that its advantage on LongBench stems from structural adaptation to very long sequences—for instance, achieving 24.9 on GovReport (3-1) (vs. 18.6 for the attention baseline) and 16.7 on HotpotQA (2-1) (vs. 13.4). Second, unlike the small gap ($\sim$2 points) between RNN and attention on short-context tasks, the difference becomes much larger on LongBench, further demonstrating the problem of full-sequence compression. We also notice that RAT(L=16) performs less favorably on code completion. However, its hybrid variant significantly improves performance in this setting, outperforming the second-best model by about four points.
Supervised Fine-tuning
Since LongBench tasks are instruction-based and pretrained-only models often struggle to follow specific prompts, we further assess the performance of SFT on some long-context QA and summarization tasks, as shown in Table 5. For NarrativeQA, where passages are extremely long and may not contain the answer within the truncated 4K-token context, we evaluate two settings: (1) summary-only, which uses only the passage summary, and (2) summary plus passage, which concatenates the summary with the full passage. In both cases, RAT(L=16)-SWA delivers strong performance while remaining efficient. Although both the summary and passage contain useful information, we observe that models, particularly the attention-based ones, perform worse in the second setting, with F1 dropping to 33. We hypothesize that in this setup, attention may be more easily distracted due to the absence of structural bias, which can lead to flattened attention score distributions [scalablesoftmax]. In contrast, models based on RAT appear more robust, likely due to their chunk-wise structure. For summarization tasks, RNN performs better than in QA, and RAT(L=16)-SWA achieves the best result on QMSum, surpassing the next-best model by approximately 1 Rouge-L point.
@@
Model Value/Slot size $^{\textbf{1}} NarrativeQA$$^{\textbf{2}}$ QMSum WikiSum RNN $ND$ 30.5 27.4 23.2 34.3 RAT(L=16) $256ND$ 60.8 43.5 23.3 36.6 Attention $4096ND$ 61.3 33.0 23.4 36.8 Attention-SWA 4096N/2D1024N/2D4096N/2D-1024N/2D 63.3 39.6 23.4 36.7 RAT(L=16)-SWA 256N/2D1024N/2D256N/2D-1024N/2D 63.2 43.7 24.6 36.7
Table 6: Retrieval ability: Accuracy performance with exact match scoring on the Needle-in-Haystack tasks with different configurations from the RULER benchmark [ruler]. We use $T=4096$, $D=2048$, and $N=24$.
Model Value/Slot size single_1 single_2 single_3 multikey_1 multikey_2 multikey_3 multiquery multivalue
RNN NDND 99.4 99.8 0.0 2.2 0.0 0.0 5.8 5.7
RAT(L=64) $64ND$ 100.0 99.8 55.6 96.4 62.2 1.0 86.3 89.1
RAT(L=16) $256ND$ 99.6 100.0 94.6 99.6 99.6 82.6 91.2 94.8
Attention $4096ND$ 100.0 100.0 100.0 99.6 99.6 95.2 98.6 99.1
Retrieval ability
In Table 6, we evaluate the three models on challenging synthetic needle-in-haystack tasks that test retrieval ability (see Subsec. C.3 for task descriptions). It can be seen that attention consistently performs well due to its full-token access. RNN handles simpler tasks like single_1 and single_2 reasonably well, but its performance drops on harder settings. RAT(L=16) matches attention closely on most tasks but struggles with UUID ones (single_3, multikey_3) due to their difficulty and evaluation metric as exact match scoring. Meanwhile, RAT(L=64) falls between RNNs and RAT(L=16), as expected given its partial access to long-range context. These results align with the model structures: attention sees all tokens directly, recurrence compresses all history into a single state, and RAT blends both, preserving chunk-level access while reducing FLOPs.